# Unraveling hydrodynamization using ultracold 1D gases

### v2

@@ IV Post-quench state modeling
Equation (8) assumes that the state before the Bragg pulse is the ground state of a homogeneous TG gas, and that the Bragg pulse is infinitely short. In our experiments and their modeling, the initial state is inhomogeneous, because of the trap needed to confine the atoms, and the pulse has a finite duration, because its power is finite. In this section, we explore theoretically the effects of inhomogeneity and finite pulse duration on the dynamics of a single TG gas.
@@
In order to isolate the effects of inhomogeneity, we study the case in which the initial state is the ground state of a homogeneous TG gas with $N=29$ particles in a 1D system with periodic boundary conditions [we set $U(z)=0$], and apply a $10$ $\mu$s Bragg scattering pulse [Eq. (7)]. We select the same number of particles and apply the same Bragg pulse as in Sec. II.2. To most cleanly compare the trapped and homogeneous cases, we choose the size of the homogeneous system so that the energy per particle before the Bragg pulse is the same as in the trapped case. The momentum and rapidity distributions obtained right after the Bragg pulse, for both the trapped and homogeneous case, are plotted in Figs. 7(a) and 7(b), respectively. Away from $p=\pm 2nk_{0}$, for which small differences are seen in Fig. 7(a), the momentum distributions of the trapped and the homogeneous systems are indistinguishable from each other. The corresponding rapidity distributions in Fig. 7(b) are also very close to each other, only differing, as expected, by the fact that in the homogeneous case there are Fermi sea-like discontinuities at the edges of each peak.
Figure 8: Time evolution of $E_{\bar{p}}$ for the homogeneous TG gas. The open symbols show the results after: (a) a 10 $\mu$s Bragg pulse, and (b) a Kapitza-Dirac pulse [as per Eq. (8)] with the same area under the pulse as in (a). The dashed lines show the results for the trapped TG gas reported in Fig. 3(b). The number of particles, $N=29$, and the energy per particle before the pulse are the same for the homogeneous TG gas as for the trapped TG gas. To obtain each curve, we integrated within a momentum interval of 0.2 $\hbar k$. The upper $x$ axis shows the dimensionless quantity $t_{\rm ev}/T^{c}_{\rm hd}$ [see Eq. (9)].
@@
To isolate the effects of the finite duration of our Bragg pulse, in Figs. 7(a) and 7(b) we also plot the momentum and rapidity distributions, respectively, for the state predicted by Eq. (8) for a Kapitza-Dirac pulse with the same area as our Bragg pulse applied to the ground state of the homogeneous case. Figure 7(b) shows that the main effect of the finite duration of the Bragg pulse is that it reduces the transfer of quasiparticles out of the initial Fermi sea, resulting in a slightly lower population of its $n=\pm 1,\pm 2$ “replicas”. Figure 7(a) shows that this results in slightly lower $n=\pm 1,\pm 2$ peaks in the momentum distribution after the Bragg pulse. Away from $p=\pm 2nk_{0}$, the momentum distributions are all indistinguishable from each other.
@@
In Fig. 8, we show the evolution of the integrated kinetic energy $E_{\bar{p}}$ after the Bragg pulse [Fig. 8(a)] and after the Kapitza-Dirac pulse [Fig. 8(b)] for the homogeneous TG gas (empty symbols). Since the two are very similar to each other, it is clear that the finite duration of the pulse, for which there can be many-body evolution during the pulse, does not in fact introduce qualitative changes. In Figs. 8(a) and 8(b) we also show, as thin dashed lines, the results for the trapped system already reported in Fig. 3(b). The evolution in the homogeneous and trapped cases are very similar. In all cases there is a rapid redistribution of kinetic energy, and oscillations of $E_{\bar{p}}$, in a time $T^{c}_{\rm hd}/2$. Hence, hydrodynamization, as seen in the experiments and in their corresponding modeling using a trapped TG gas, is qualitatively affected by neither the fact that the system is inhomogeneous nor by the finite duration of the Bragg pulse. Therefore, modeling that directly uses Eq. (8), with the appropriate initial state and area of the Bragg pulse, provides a quantitatively reliable description of the dynamics of the trapped TG gas following the finite-duration Bragg pulse.
@@
Having shown that hydrodynamization is essentially the same in the presence and absence of a trap, as well as with and without finite-time duration pulses, we next explore the local prethermalization of the momentum distribution in the presence and absence of a trap. Local prethermalization refers to the equilibration of the local momentum distribution to the generalized Gibbs ensemble (GGE) prediction [25, 26, 27, 28]. During local prethermalization, the overall density distribution does not change significantly. After local prethermalization, trapped systems exhibit global dynamics during which the density distribution changes and with it the local and global rapidity and momentum distributions. Such a long-wavelength and long-time dynamics in an integrable system is described by generalized hydrodynamics (GHD) [29, 30, 31, 32, 33], as shown in recent experiments [34, 35, 36, 37, 38].
@@
Figure 9: Time evolution of the momentum distributions of trapped and homogeneous TG gases. Time evolution of $\Delta f$ [see Eq. (13)] for a trapped gas after a 10 $\mu$s Bragg pulse (circles), a homogeneous gas after a 10 $\mu$s Bragg pulse (squares), and a homogeneous gas after a Kapitza-Dirac pulse [as per Eq. (8)] with the same area under the pulse as the other two cases (triangles). In all three cases, the number of particles is $N=29$, and the energy per particle is the same before the pulse. (Inset) Time evolution of $f(p=0)$ for the same systems considered in the main panel. The vertical grey band highlights the time interval used to calculate the time-averaged $\bar{f}(p)$ input in the calculation of $\Delta f(t_{\rm ev})$.
@@
The local prethermalization time for $f(p)$ is in general momentum-dependent, with the occupation of lower momenta (determined by correlations over longer distances) requiring longer prethermalization times [see Fig. 1(b)]. We will discuss this dependence in more detail in the next section. In the inset in Fig. 9, we plot the time evolution of $f(p=0)$, which requires the longest time to prethermalize. We show results following the Bragg pulse in the trapped and homogeneous systems and after the Kapitza-Dirac pulse in the homogeneous system. They all behave qualitatively similarly except at long times. The inset in Fig. 9 shows that after equilibrating to a nearly time-independent value, $f(p=0)$ in the trapped system departs from that value at $t\sim 1$ ms. This departure occurs because the density distribution begins to change in the trap. In the homogeneous cases the density distribution does not change and $f(p=0)$ remains unchanged (except for the inevitable revivals that occur in finite systems).
In the inset in Fig. 9 one can see that the values of $f(p=0)$ for the three systems considered are nearly time independent in the interval $t_{\rm ev}\in[200,300]\,\mu$s highlighted by the grey band. We use 6 times spaced by $\delta t_{\rm ev}=20\mu s$ in that interval to calculate the time-averaged momentum distribution $\bar{f}(p)$. We take $\bar{f}(p)$ to be the locally equilibrated momentum distribution. To characterize the local prethermalization of the momentum distribution, we then calculate the relative difference
[…]
@@
The main panel in Fig. 9 shows the time evolution of $\Delta f$ for the trapped TG gas after a 10 $\mu$s Bragg pulse, and for the homogeneous TG gas after a 10 $\mu$s Bragg pulse and after a Kapitza-Dirac pulse. The three curves are very similar to each other, differing only at the earliest time, presumably because the Bragg pulse calculations get a local prethermalization head start during the pulse. The delay seems unimportant for the characterization of the local prethermalization of $f(p)$. As with hydrodynamization, local prethermalization is essentially the same in the presence and absence of a trap, as well as with and without pulses with a finite-time duration.
V Hydrodynamization and the lack thereof in finite superpositions of plane waves
@@
Having shown that there is no need to introduce the trapping potential or to carry out a finite-duration pulse to reproduce the experimental observations, we next take a final step in simplifying our theoretical modeling. In Eq. (8), the quasiparticles are in infinite superpositions of plane waves, corresponding to the infinite Bragg peaks whose weights are determined by the area under the pulse via the Bessel function. In this section, we consider the cases in which only the three or two dominant Bragg peaks from the experiments in Secs. II.2 and III.2, respectively, are present. We study their time evolution under a translationally invariant Hamiltonian.
For all the calculations in this section, we fix the length of our periodic 1D system to be $L=20\times 2\pi/k_{0}$, where $k_{0}$ is the wavenumber from the experiments discussed in previous sections. To study the effect that changing the particle density (the Fermi momentum $\theta_{F}$) has on the different time scales, we change the particle number $N$.
V.1 Three-peak states
We first consider states that have a three-modal rapidity distribution, which we refer to as three-peak states. They have wave functions of the form
where we have chosen the phases of each “replica” of the initial Fermi sea (with Fermi momentum $\theta_{F}$) to match those of the Bessel functions in Eq. (8). Choosing other phases introduces phase-shifts with respect to the results in the previous sections, but does not qualitatively change any of the scalings discussed in what follows. We select the weights of the peaks to be $A_{0}=2/\sqrt{6}$ and $A_{\pm 1}=1/\sqrt{6}$, which are different from the values $A_{0}\approx\sqrt{2}/2$ and $A_{\pm 1}\approx 1/2$ that we used in Sec. IV when modeling the experiment. As we will see, changing $A_{0}$ and $A_{\pm 1}$ does not qualitatively change the hydrodynamization and prethermalization phenomenology.
In Fig. 10, we plot the time evolution of $E_{\bar{p}}$ in a homogeneous system with $N=33$ bosons with periodic boundary conditions for the three-peak state, showing results for $k=3k_{0}$ [Fig. 10(a)] and $k=5k_{0}$ [Fig. 10(b)]. When plotted in terms of the dimensionless time $t_{\rm ev}/T^{c}_{\rm hd}$ (see the upper $x$ axes labels), the energy curves for different $k$ exhibit qualitatively similar short-time behavior. For example, the energies $E_{\bar{p}}$ for $\bar{p}\sim\hbar k$ initially increase rapidly, and exhibit one period of oscillation, within $t_{\rm ev}\approx T^{c}_{\rm hd}/2$. Those results are qualitatively similar to the ones reported in Fig. 8 although, because $T_{\rm dp}/T^{c}_{\rm hd}=\hbar k/(2\theta_{F})$ is greater, more hydrodynamization oscillations occur in Fig. 10 for $p\approx\hbar k$. The hydrodynamization oscillation period does not significantly depend on $p$, but the hydrodynamization damping times and the local prethermalization times do depend on $p$.
Next, we study the prethermalization of the momentum distribution in more detail. As shown in Refs. [25, 39, 40], the integrability of the Tonks-Girardeau gas implies that the momentum distribution after equilibration is not thermal. Instead, it is described by a GGE that is fully determined by the conserved rapidity distribution of the underlying quasiparticles. Similar results have been obtained in other paradigmatic integrable models that map onto quadratic ones, such as the Luttinger model [41] and the transverse-field Ising model in 1D [27, 42, 43]. In Fig. 11, we show $f(p)$ for the state studied in Fig. 10(a) at $t_{\rm ev}=0$ and 40 $\mu$s, along with the GGE prediction. Except for the occupation of $p=0$, which requires a longer time to equilibrate, $f(p)$ at $t_{\rm ev}=40$ $\mu$s is indistinguishable from the GGE prediction.
Figure 12: Time evolution of $f(p)$ within the central peak for three-peak states. The initial states have: (a) $N=33$ ($\theta_{F}=0.825\hbar k_{0}$) and $k=3k_{0}$, (b) $N=33$ ($\theta_{F}=0.825\hbar k_{0}$) and $k=5k_{0}$, and (c) $N=65$ ($\theta_{F}=1.625\hbar k_{0}$) and $k=3k_{0}$. Note the change in the times shown in each panel. The upper $x$ axis shows the dimensionless quantity $t_{\rm ev}/T_{\rm dp}$. The legends show the value of $p/\theta_{\rm F}$. The vertical grey bands ($t_{\rm ev}\in[0,1]T_{\rm dp}$) highlight the time scale in which the hydrodynamization oscillations damp out.
How the occupation $f(p)$ of each momentum mode approaches the GGE prediction depends in a complicated way on $p$, $k$, and $\theta_{F}$. In Fig. 12, we show the time evolution of $f(p)$ for $p<2\theta_{F}$ (i.e., within the central momentum peak), for $N=33$ when $k=3k_{0}$ [Fig. 12(a)] and when $k=5k_{0}$ [Fig. 12(b)], and for $N=65$ when $k=3k_{0}$ [Fig. 12(c)]. Note that the absolute time axes differ across the panels so that the dimensionless (upper) time axes $t_{\rm ev}/T_{\rm dp}$ [see Eq. (18)] are the same. The results in Fig. 12 show that, for any given pair of values of $k$ and $\theta_{F}$, the shapes of the $f(p)$ vs $t_{\rm ev}$ curves change when changing $p$, with the prethermalization time increasing as $p$ approaches 0. Fixing $p/\theta_{F}$ (which have the values reported in the legends) and changing $k$ or $N$ also results in changes in the $f(p)$ curves. There are, however, two universal features that can be seen across Figs. 12(a)–12(c). The first is that the damping of the hydrodynamization oscillations in all panels and for all $p$ occurs within $t_{\rm ev}\simeq T_{\rm dp}$ (highlighted by the shaded bands). The second is that the prethermalization time for $f(p)$, i.e., the time after which the $f(p)$ curves become nearly time independent, are momentum independent when $p\gtrsim\theta_{F}$.
The diversity of shapes among the $f(p)$ curves in Fig. 12 make it clear that defining the prethermalization time for different momentum modes $p$ is a challenging task, even if one were to smooth out the hydrodynamization oscillation. Addressing this challenge is postponed to future work. The results in Fig. 12 also hint that the prethermalization times for all momentum modes are proportional to $T_{\rm dp}$, because one can see that for each value of $p/\theta_{F}$, equilibration in different panels occurs at about the same value of $t_{\rm ev}/T_{\rm dp}$. Building on this observation, we next show that the prethermalization time scale for the entire momentum distribution is indeed proportional to $T_{\rm dp}$.
In Fig. 13, we plot the time evolution of $\Delta f$ [see Eq. (13)] vs $t_{\rm ev}/T_{\rm dp}$ for systems with the same number of particles ($N=11$) and different $k$ [Fig. 13(a)], and for systems with the same value of $k=3k_{0}$ and different $N$ [Fig. 13(b)]. In both panels, one can see that the curves of $\Delta f$ vs $t_{\rm ev}/T_{\rm dp}$ collapse onto each other, except for the $N=11$ curve past $t_{\rm ev}\sim 1.5T_{\rm dp}$ in Fig. 13(b), which deviates due to finite-size effects. For times $0.5T_{\rm dp}\lesssim t_{\rm ev}\lesssim 3T_{\rm dp}$, the results for $N>11$ are well described by an exponential function. Specifically, $\Delta f(t)\propto\exp(-b\,t_{\rm ev}/T_{\rm dp})$, which means that the prethermalization time $T_{\rm lp}$ for the momentum distribution is proportional to $T_{\rm dp}$. [Once $\Delta f(t)$ becomes a fraction of a percent near $t_{\rm ev}\sim 3T_{\rm dp}$, the evolution slows down, transitioning to a roughly polynomial decay in time (see the inset in Fig. 13(b). Since the effect may be too small to experimentally detect, we will not explore it further here.] It has previously been argued that the quantum dynamics of TG gas momentum distributions can be like those of generic observables in interacting integrable models that cannot be mapped onto noninteracting models [44]. We therefore expect our TG results to qualitatively describe dynamics away from the TG limit, where the Lieb-Liniger model cannot be mapped onto a noninteracting model, with a proportionality constant between $T_{\rm lp}$ and $T_{\rm dp}$ that depends on the contact interaction strength $g$.
In contrast, we do not expect the TG gas to serve as a model for the quantum dynamics of the density distribution in the experiments. That is because the matrix elements of few-body observables in noninteracting models are qualitatively different from those in interacting models [44]. The density dynamics of a TG gas and a noninteracting Fermi gas are identical, while there is no such mapping for a Lieb-Liniger gas away from the TG limit. The Gaussian nature of the energy eigenstates in noninteracting models makes the observables equilibrate polynomially in time [45, 46, 47, 48, 49], as opposed to exponentially.
@@ V.1 Three-peak states
vs $t_{\rm ev}/T_{\rm dp}$. Results for $\Delta\rho(t_{\rm ev})$ are shown for systems with the same number of particles ($N=11$) and different $k$ [Fig. 14(a)], and for systems with the same value of $k=3k_{0}$ and different $N$ [Fig. 14(b)], which parallel the results for $\Delta f(t_{\rm ev})$ in Figs. 13(a) and 13(b), respectively. In contrast to $\Delta f(t_{\rm ev})$, $\Delta\rho(t_{\rm ev})$ exhibits large fast changes on the hydrodynamization time scale, the result of the oscillations of $\rho(z,t_{\rm ev})$ discussed earlier (and shown in Fig. 1c of Ref. [5]). There are also oscillations with the longer $T_{\rm dp}$ period. Most notably and as expected, the amplitude of the oscillations decays polynomially (as opposed to exponentially) in time as $(t_{\rm ev}/T_{\rm dp})^{-1}$, as demarcated by the dashed lines in the plots. We expect this decay to change into an exponential decay as one moves away from the TG limit, a change that could be explored in future studies.
We close our discussion of the three-peak case by connecting our results for prethermalization here to the local prethermalization results reported in Ref. [5]. There we studied the time evolution of “$p_{\rm f}$”, which is defined so that ±$\pm p_{\rm f}$ are the momenta enclosing $f\%$ of the atoms in the central peak. Since the results were found to be qualitatively similar for $f$ in the range of 40 to 60, we focused on $f=50$,

### v3

@@ Appendix C Insensitivity to the trap and
Equation (17) assumes that the state before the Bragg pulse is the ground state of a homogeneous TG gas, and that the Bragg pulse is infinitely short. In our experiments and their modeling, the initial state is inhomogeneous, because of the trap needed to confine the atoms, and the pulse has a finite duration, because its power is finite. In this appendix, we explore theoretically the effects of inhomogeneity and the Bragg pulse used on the dynamics of a single TG gas.
In order to isolate the effects of inhomogeneity, we study the case in which the initial state is the ground state of a homogeneous TG gas with $N=29$ particles in a 1D system with periodic boundary conditions [we set $U(z)=0$], and apply a $10$ $\mu$s Bragg scattering pulse [Eq. (16)]. We select the same number of particles and apply the same Bragg pulse as in Appendix B. To most cleanly compare the trapped and homogeneous cases, we choose the size of the homogeneous system so that the energy per particle before the Bragg pulse is the same as in the trapped case. The momentum and rapidity distributions obtained right after the Bragg pulse, for both the trapped and homogeneous case, are plotted in Figs. 14(a) and 14(b), respectively. Away from $p=\pm 2nk_{0}$, for which small differences are seen in Fig. 14(a), the momentum distributions of the trapped and the homogeneous systems are indistinguishable from each other. The corresponding rapidity distributions in Fig. 14(b) are also very close to each other, only differing, as expected, by the fact that in the homogeneous case there are Fermi sea-like discontinuities at the edges of each peak.
@@
To isolate the effects of the finite duration of our Bragg pulse, in Figs. 14(a) and 14(b) we also plot the momentum and rapidity distributions, respectively, for the state predicted by Eq. (17) for a Kapitza-Dirac pulse with the same area as our Bragg pulse applied to the ground state of the homogeneous case. Figure 14(b) shows that the main effect of the finite duration of the Bragg pulse is that it reduces the transfer of quasiparticles out of the initial Fermi sea, resulting in a slightly lower population of its $n=\pm 1,\pm 2$ “replicas”. Figure 14(a) shows that this results in slightly lower $n=\pm 1,\pm 2$ peaks in the momentum distribution after the Bragg pulse. Away from $p=\pm 2nk_{0}$, the momentum distributions are all indistinguishable from each other.
@@
In Fig. 15, we show the evolution of the integrated kinetic energy $E_{\bar{p}}$ after the Bragg pulse [Fig. 15(a)] and after the Kapitza-Dirac pulse [Fig. 15(b)] for the homogeneous TG gas (empty symbols). Since the two are very similar to each other, it is clear that the finite duration of the pulse, for which there can be many-body evolution during the pulse, does not in fact introduce qualitative changes. In Figs. 15(a) and 15(b) we also show, as thin dashed lines, the results for the trapped system already reported in Fig. 13(b). The evolution in the homogeneous and trapped cases are very similar. In all cases there is a rapid redistribution of kinetic energy, and oscillations of $E_{\bar{p}}$, in a time $T^{c}_{\rm hd}/2$. Hence, hydrodynamization, as seen in the experiments and in their corresponding modeling using a trapped TG gas, is not qualitatively affected by either the fact that the system is inhomogeneous or by the finite duration of the Bragg pulse. Therefore, modeling that directly uses Eq. (17), with the appropriate initial state and area of the Bragg pulse, provides a quantitatively reliable description of the dynamics of the trapped TG gas following the finite-duration Bragg pulse.
Figure 15: Time evolution of $E_{\bar{p}}$ for the homogeneous TG gas. The open symbols show the results after: (a) a 10 $\mu$s Bragg pulse, and (b) a Kapitza-Dirac pulse [as per Eq. (17)] with the same area under the pulse as in (a). The dashed lines show the results for the trapped TG gas reported in Fig. 13(b). The number of particles, $N=29$, and the energy per particle before the pulse are the same for the homogeneous TG gas as for the trapped TG gas. To obtain each curve, we integrated within a momentum interval of 0.2 $\hbar k_{0}$. The upper $x$ axis shows the dimensionless quantity $t_{\rm ev}/T^{c}_{\rm hd}$ [see Eq. (2)].
@@
Next, we explore the local prethermalization of the momentum distribution in the presence and absence of a trap. Local prethermalization refers to the equilibration of the local momentum distribution to the GGE prediction [20, 27, 24, 28]. During local prethermalization, the overall density distribution does not change significantly. After local prethermalization, trapped systems exhibit global dynamics during which the density distribution changes and with it the local and global rapidity and momentum distributions. Such a long-wavelength and long-time dynamics in an integrable system is described by generalized hydrodynamics [9, 10, 11, 12, 13], as shown in recent experiments [52, 16, 53, 54, 18].
@@
Figure 16: Time evolution of the momentum distributions of trapped and homogeneous TG gases. Time evolution of $\Delta f$ [see Eq. (5)] for a trapped gas after a 10 $\mu$s Bragg pulse (circles), a homogeneous gas after a 10 $\mu$s Bragg pulse (squares), and a homogeneous gas after a Kapitza-Dirac pulse [as per Eq. (17)] with the same area under the pulse as the other two cases (triangles). In all three cases, the number of particles is $N=29$, and the energy per particle is the same before the pulse. (Insets) Time evolution of $f(p=0)$ (upper inset) and $f(p=2\hbar k_{0})$ (lower inset) for the same systems considered in the main panel. The vertical gray bands highlight the time interval used to calculate the time-averaged $\bar{f}(p)$ input in the calculation of $\Delta f(t_{\rm ev})$.
@@
The local prethermalization time for $f(p)$ is in general momentum-dependent, with the occupation of lower momenta (determined by correlations over longer distances) requiring longer prethermalization times. We discuss this dependence in detail in Sec. III. The insets in Fig. 16 show the time evolution of $f(p=0)$ (top inset) and $f(p=2\hbar k_{0})$ (bottom inset), both following the Bragg pulse in the trapped (red points) and homogeneous systems (green points) and after the Kapitza-Dirac pulse in the homogeneous system (blue points). The initial changes are qualitatively the same for the different pulses. As expected, $f(p=0)$ takes longer to prethermalize. After first equilibrating to a nearly time-independent value, $f(p=0)$ and $f(p=2\hbar k_{0})$ in the trapped system depart from that value at $t\sim 1$ ms because the density distribution begins to change in the trap. In the homogeneous cases the density distribution does not change and $f(p=0)$ remains unchanged (except for the inevitable revivals that ultimately occur in finite systems).
@@
In the insets in Fig. 16 one can see that the values of $f(p=0)$ and $f(p=2\hbar k_{0})$ for the three systems considered are nearly time-independent in the interval $t_{\rm ev}\in[200,300]\,\mu$s highlighted by the gray band. We use six times spaced by $\delta t_{\rm ev}=20\mu s$ in that interval to calculate the time-averaged momentum distribution $\bar{f}(p)$. We take $\bar{f}(p)$ to be the locally equilibrated momentum distribution. To characterize the local prethermalization of the momentum distribution, we calculate the relative difference $\Delta f(t_{\rm ev})$ [see Eq. (5)]. The main panel in Fig. 16 shows the time evolution of $\Delta f$ for the trapped TG gas after a 10 $\mu$s Bragg pulse, and for the homogeneous TG gas after a 10 $\mu$s Bragg pulse and after a Kapitza-Dirac pulse. The three curves are very similar to each other, differing only at the earliest time, presumably because the Bragg pulse calculations get a local prethermalization head start during the pulse. The delay seems unimportant for the characterization of the local prethermalization of $f(p)$. As with hydrodynamization, local prethermalization is essentially the same in the presence and absence of a trap, as well as with and without pulses with a finite-time duration.
@@
Having shown that there is no need to introduce the trapping potential or to carry out a finite-duration pulse to reproduce the experimental observations, in the main text we take a final step in simplifying our theoretical modeling. In Eq. (17), the quasiparticles are in infinite superpositions of plane waves, corresponding to the infinite Bragg peaks whose weights are determined by the area under the pulse via the Bessel function. In the main text, we consider the cases in which only three or two Bragg peaks are present, to avoid the small complications associated with the small higher order peaks.
Appendix D Three-peak states:
Density distribution and $p_{50}$
@@ Appendix D Three-peak states:
In contrast to the momentum distribution, we do not expect the TG gas to serve as a model for the quantum dynamics of the density distribution in the experiments. That is because the matrix elements of few-body observables in noninteracting models are qualitatively different from those in interacting models [29]. The density dynamics of a TG gas and a noninteracting Fermi gas are identical, while there is no such mapping for a Lieb-Liniger gas away from the TG limit. The Gaussian nature of the energy eigenstates in noninteracting models makes the observables equilibrate polynomially in time [55, 56, 57, 58, 59], as opposed to exponentially.
@@
vs $t_{\rm ev}/T_{\rm dp}$. Results for $\Delta\rho(t_{\rm ev})$ are shown for systems with the same number of particles ($N=11$) and different $k$ [Fig. 17(a)], and for systems with the same value of $k=3k_{0}$ and different $N$ [Fig. 17(b)], which parallel the results for $\Delta f(t_{\rm ev})$ in Figs. 5(a) and 5(b), respectively. In contrast to $\Delta f(t_{\rm ev})$, $\Delta\rho(t_{\rm ev})$ exhibits large fast changes on the hydrodynamization timescale, the result of the oscillations of $\rho(z,t_{\rm ev})$ discussed earlier (and shown in Fig. 1(c) of Ref. [8]). There are also oscillations with the longer $T_{\rm dp}$ period. Most notably and as expected, the amplitude of the oscillations decays polynomially (as opposed to exponentially) in time as $(t_{\rm ev}/T_{\rm dp})^{-1}$, as demarcated by the dashed lines in the plots (see also Ref. [60]). We expect this decay to change into an exponential decay as one moves away from the TG limit, a change that could be explored in future studies.
We close our discussion of the three-peak case by connecting our results for prethermalization in this work to the local prethermalization results reported in Ref. [8]. There we studied the time evolution of “$p_{\rm f}$,” which is defined so that ±$\pm p_{\rm f}$ are the momenta enclosing $f\%$ of the atoms in the central peak. Since the results were found to be qualitatively similar for $f$ in the range of 40 to 60, we focused on $f=50$,